Shift report 20030320 08:00-16:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Thu Mar 20 2003 - 15:50:45 EST

  • Next message: Flemming Videbaek: "remaining dAu Run Plan"
    ------------------------------------------------------------
    Date: Thu Mar 20 15:50:44 2003, Shift: 08:00-16:00
    Supervisor: SL, with: 
    Log book#14 pages 68 - 68
    
    Shift Summary:
    ------------------------------------------------------------
    1.Good beam since 9am, zdc at 10000, tofw 30000
    Store ended around 1430 after voting with phenix for dump. ZDC below 3000.
    2.Runs 8525 through 8528 with fs at 4 deg 1/1B and mrs at 40 deg 2000B, altogether approx 60k tr 2 events. Runs 8529 through 8532 test and pedestal runs
    3.T3 off, T4 and T5 tripped just after nine, but ok rest of the day.
    ------------------------------------------------------------
    


    This archive was generated by hypermail 2.1.5 : Thu Mar 20 2003 - 15:55:53 EST